What is the concentration of hydroxide ions in 750 mL of a solution containing the maximum amount of dissolved calcium hydroxide?


1
Expert's answer
2022-03-29T16:39:03-0400

Solubility of calcium hydroxide (Ca(OH)2) = 1.73 g/L;

In 750 mL (0.750 L): m(Ca(OH)2) = 1.73 * 0.75 = 1.30 g;

M(Ca(OH)2) = 74 g/mol:

n(Ca(OH)2) = m(Ca(OH)2)/M(Ca(OH)2) = 1.30/74 = 0.018 mol;

n(OH-) = 2 * n(Ca(OH)2) = 2 * 0.018 mol = 0.036 mol;

C(OH-) = n(OH-)/0.750 = 0.036/0.75 = 0.05 mol/L.

Answer: 0.05 mol/L.